OpenStudy (donnie1999):

What is the inverse of the function below? f (x) = 2x + 12 _________ 5

OpenStudy (donnie1999):

@dcskaf

OpenStudy (anonymous):

x= 2y+12 ----- 5 multiply 5 then subtract and finally divide

OpenStudy (donnie1999):

i got f-^1(X)= 14/5 x ???

OpenStudy (anonymous):

Ok lets go through it together x= 2y+12 ----- 5 multiply 5 you get 5x=2y+12

OpenStudy (anonymous):

ok so far?

OpenStudy (donnie1999):

ya

OpenStudy (anonymous):

then you subtract 12 to get 5x-12 =2y

OpenStudy (donnie1999):

so would it be f-^1(x)= -2x+12 ------ ???? 5

OpenStudy (anonymous):

not quite it should be y= 5x-12 ----- 2 which simplifies to either 2.5x-6 or 5/2x-6

OpenStudy (anonymous):

do you get it?

OpenStudy (donnie1999):

idk would it be f-^1(x)= 5/2x - 6 then???

OpenStudy (anonymous):

yes